A uniform brick of length 20 m is placed
over the edge of a horizontal surface with a
maximum overhang of 10 m attained without
tipping.Now two identical uniform bricks of length
20 m are stacked over the edge of a horizontal
surface.What maximum overhang is possible for
the two bricks (without tipping)?
Answer in units of m. Don't round answer.
1
Expert's answer
2020-01-21T03:32:31-0500

The center of first rod will lie exactly at the edge of first rod and then the center of mass of two combined rods will lie at the edge of the table.

So now the whole system will rest on the rod and it will not tipping off.

Since both rods are identical so we can say that the system will have its center of mass at the mid point on the line joining the two centers

So the value of x will be at mid point of line joining the two points on rod


"x=0.5(10)=5\ m"

So total length over the edge will be given as


"L=10+5=15\ m"